4. A girl weighing 50kg is in a 70kg cart, resting. The girl jumps out of the carriage from her back at a speed of 5m / s. Find the size and speed of the cart after the girl is thrown. Friction is not taken into account.


1
Expert's answer
2021-02-15T00:58:17-0500

From the law of conservation of momentum, we have:


"p_i=p_f,""m_{girl}v_{girl}+m_{cart}v_{cart}=0,""v_{cart}=-\dfrac{m_{girl}v_{girl}}{m_{cart}},""v_{cart}=-\dfrac{50\ kg\cdot5\ \dfrac{m}{s}}{70\ kg}=-3.57\ \dfrac{m}{s}."

The sign minus means that the cart moves in the opposite direction to the motion of the girl.